OpenStudy (anonymous):

6x^x=-12x-2 using quadratic formula

OpenStudy (anonymous):

i got \[\frac{ -12\pm \sqrt{96} }{ 12 }\] any way of simplifying it? i do know more or less with radicals but i cant with this huge number

OpenStudy (luckycoins888):

i dont know

OpenStudy (luckycoins888):

@crazychickengirl123 really?

OpenStudy (anonymous):

@mustard84 your working out was right you can simplify \[\sqrt{96}\] to \[\sqrt{16\times6}\] so that will be \[\sqrt{16}\times \sqrt{6}\] which give you \[4\sqrt{6}\]

OpenStudy (anonymous):

also using a calculator can give you the exact answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no worries..